OpenStudy (anonymous):

Given a = 2, b = -2, and k = 3, what is the equation of the graph if the parent function is y equals the square root of x? A. y equals two times the square root of 2x, plus 3 B. y equals negative two times the square root of 2x, plus 3 C. y equals two times the square root of negative 2x, minus 3 D. y equals two times the square root of negative 2x, plus 3

OpenStudy (anonymous):

What material is this from?

OpenStudy (anonymous):

Algebra 2/trig

OpenStudy (anonymous):

In their notation is "a" the horizontal shift, "b" the vertical shift, and "k" the vertical scaling?

OpenStudy (anonymous):

i have no idea.. i have to do it online and i couldn't understand the material very well and i meet with my math tutor next wednesday but i have to have it done..

OpenStudy (anonymous):

So for example the parent function y=x y=k(x-b)+a? The problem is those variable names are arbitrary it could also be y=a(x-k)+b etc. Is there an example in your text that shows their usage?

OpenStudy (anonymous):

nope its all online.. thats all that it gave me

ganeshie8 (ganeshie8):

check in ur online lessons page, if u have some equation like below :- \(\large f(x) = b\sqrt{ax} + k\)

ganeshie8 (ganeshie8):

which online school u doing ? flvs/vlacs/connexus/others ?

OpenStudy (anonymous):

\[y=2\sqrt{2x+3}\] or\[y=2\sqrt{2x}+3\]

OpenStudy (anonymous):

I'm doing advanced academics

OpenStudy (anonymous):

ok so what would this be? For the parent function y equals the cube root of x, what effect does the values h = -4 have on the graph? A. Horizontal shift of 4 units to the right B. Horizontal shift of 4 units to the left C. Vertical shift of 4 units up D. Vertical shift of 4 units down

OpenStudy (anonymous):

I don't know the conventions that your course uses for "h". I can't believe that it wouldn't be listed for you somewhere in your material.

OpenStudy (anonymous):

it would be D right?

ganeshie8 (ganeshie8):

h usually refers to horizontal shift. again we need to know wat they refer to as h, it will be there in ur lessons page im sure :)

OpenStudy (anonymous):

yayyy thanks! i got a 70% but i got all the ones right that you guys gave me!! thank you!!!!!

ganeshie8 (ganeshie8):

:/

ganeshie8 (ganeshie8):

70% is not bad, congrats :))

ganeshie8 (ganeshie8):

however dont say it loud lol, helping others in test/quiz is a crime in openstudy :|